# Death by Impact: Evidence for Merger-Driven Quenching in a Collisional Ring Galaxy at Cosmic Noon

### v1

@@ 4.2 Stellar Kinematics
We measure the stellar kinematics in each region using pPXF (Cappellari, 2017) to model the absorption lines in the Y-band spectra. The stellar continuum is modeled with 32 simple stellar population templates from the EMILES library (Vazdekis et al., 2016). These templates are based on the Chabrier (2003) initial mass function, BaSTI isochrones (Pietrinferni et al., 2004), and include eight stellar ages spanning from 0.15 to 14 Gyr, logarithmically spaced in steps of 0.22 dex. Four metallicity values were considered: [Z/H] = –1.5, –0.35, 0.06, and 0.4. During the fitting process, we applied multiplicative polynomials of sixth order to account for variations in the spectral shape. Note that the H emission line was fit with a velocity allowed to vary within km s-1 relative to the stellar component, to accommodate potential kinematic offsets.
@@
In conclusion, analysis of the kinematics shows that the Bullet galaxy is gas-poor and is moving toward the observer, while the Upper Ring is gas-rich and moving away from the observer. While the Bullet and the Upper Ring appear connected in the imaging, this must be clearly a projection effect, with the Bullet likely being in the foreground of the Upper Ring.
@@ 4.3 Prospector Fit and SFHs
Similarly to what done with the photometry and explained in Sec. 3, we use prospector to model the SED of different regions, but this time including both photometric data and Y-band spectra. The model includes 29 free parameters that describe the stars, gas, and dust, as explained earlier, with additional parameters necessary to fit the spectroscopy, accounting for example for imperfect flux calibration and spectral uncertainties, as explained in detail by Park et al. (2024). It is important to note that while dust emission is considered in our models, AGN and nebular emission are not included in our current model. For this reason, we mask the H emission lines in the spectra and do not include the H-band spectra in the fit.
The best-fit Prospector model and the observed data are presented in Fig. 4 for the Bullet and Host galaxies. We fit MOSFIRE Y-band spectra in combination with JWST and HST photometric data, extracted from the same spatial regions. Notably, the Host spectrum appears to exhibit a small excess absorption at the Ca K line (CaII Å), which is not fully accounted for by the stellar population template. This excess is also visible in Fig. 3, where the Y band spectrum is modeled using pPXF. The Ca K line has attracted growing attention in recent literature (Belli et al., 2024; Liboni et al., 2025) because it is a resonant transition and thus can be used to detect cold gas in the galaxy reservoir and/or outflows, as an alternative diagnostic to the Na D line. The gas traced is mostly neutral, given that the ionization potential of Ca II (11.9 eV) is lower than that of hydrogen.
In line with the objectives of this study, Fig. 5 presents the SFHs of the Bullet and Host derived using two methods: (1) combined photometry and spectroscopy, and (2) photometry alone. In both cases, the photometric data were extracted from the same spatial regions, as defined by the boxes shown in Fig. 3. The results demonstrate a high degree of consistency between the two approaches, with both yielding comparable present-day SFRs (first age bin) and similar burst/quenching patterns. Noticeable discrepancy appears in the older age bins of the Host galaxy’s SFH. However, this variations do not affect the overall interpretation or conclusions of this work. Note that the associated uncertainties of the derived SFR in each age bin represent the confidence intervals around the maximum a posteriori estimates.
@@ 5.1 Quenching During the Impact Window
The kinematics of the collisional ring enables a reconstruction of the time elapsed since the impact between the two galaxies (see Sec. 4.1). We assume that the ring started expanding at the time of impact, and that the expansion rate was constant (see also Renaud et al. 2018 and references therein). Accounting for observational uncertainties, the estimated collision time ranges from 47 to 96 Myr ago. This measurement is consistent with the ring detectability range ( Myr, Renaud et al., 2018; Mart´ınez-Delgado et al., 2023) and observed rings lifetime in previous studies at lower redshifts ( 50 to 300 Myr, Conn et al., 2011; Ditrani et al., 2024; Smith et al., 2012). The 47 Myr lower bound corresponds to the maximum inclination and minimum rotational velocity. Recent studies (e.g., Scudder et al., 2012; Patton et al., 2013; Sparre & Springel, 2016; Renaud et al., 2018) suggest that changes in the SFH of interacting galaxies can begin well before first contact or final coalescence (at separations kpc), sometimes as early as Myr prior to impact. Incorporating this pre-collision influence, we estimate a window of approximately 50-150 Myr over which the interaction can impact the SFH.
@@
The timeline of the impact is compared to the SFH of the Host and the Bullet in Fig. 6. As illustrated by the shaded regions in the figure, the impact window timeline (50–150 Myr) aligns closely with the onset of major changes in star formation activity—specifically, a pronounced starburst in the Host galaxy and rapid quenching in the Bullet. This temporal correspondence highlights the pivotal role of the merger in driving these evolutionary transitions. Quantitatively, the Host exhibits a star formation rate more than 20 times higher than at the onset of the impact window, while the Bullet’s star formation rate declines to nearly one-tenth of its pre-impact value, demonstrating contrasting responses by the two galaxies to the same interaction event.
@@
The alignment between the timeline of quenching and that of the galaxy interaction strongly suggests the presence of a physical connection between these two processes. However, in principle it is possible that this is merely a coincidence, and the Bullet was instead quenched by an unrelated mechanism while it was also interacting with the Host. The likelihood of a massive galaxy undergoing quenching independently of an external impact can be roughly estimated. Park et al. (2023) provide number density estimates for recently quenched galaxies at redshift with stellar masses comparable to our system. They report that the number density of galaxies that quenched rapidly within the past 300 Myr at this redshift is approximately . According to Weaver et al. (2023), based on COSMOS2020 data, the total number density of galaxies with comparable stellar mass at is about , with quiescent galaxies comprising roughly of that population. This means that if we select a massive quiescent galaxy at that is in a major merger, the probability that this system is undergoing quenching within this short timescale purely by chance—i.e., independently of the interaction with the companion—is of the order of . This low probability strengthens the case for merger-induced quenching in our system.
@@ 5.2 Alternative Process Associated with Galaxy Mergers
Quenching in merging systems is typically believed to occur due to the alteration of gas inflow, typically by rapidly consuming the gas reservoir of the galaxy during significant bursts of star formation (e.g., Ellison et al., 2022; Li et al., 2023). However, we rule out the possibility that quenching in the Bullet is caused by gas exhaustion based on the absence of any detectable starburst signature in the SFH (see Figs. 2 and 5). This conclusion holds even when adopting a bursty SFH model in Prospector (see Sec. 3). Another possibility is that the interaction triggers an episode of strong AGN feedback in the Bullet, which may rapidly remove the ISM gas content and shut down the star formation. In this classic scenario (e.g., Di Matteo et al., 2005; Hopkins et al., 2008), the black hole accretion is also typically accompanied by a substantial starburst, which we do not detect. In contrast, the Host galaxy exhibits both a significant starburst and AGN activity, indicating substantial gas inflow during the collision event. This suggests a gas-rich environment, where gas is funneled into the central regions as a result of the impact.
Several studies have proposed that galaxy interactions can induce significant turbulence within the ISM, which in turn may suppress star formation efficiency by disrupting the gravitational collapse of gas clouds (e.g., Ellison et al., 2018; van de Voort et al., 2018; Suess et al., 2025). This phenomenon is sometimes associated with so-called violent disc instability, a process that can be triggered during major mergers (e.g., Dekel & Burkert, 2014; Zolotov et al., 2015). In the context of our system, such a scenario remains plausible, particularly if the timescale for gas inflow is shorter than the timescale required for star formation. This condition corresponds to a wetness parameter333This parameter is defined as the ratio of the timescale required for star formation over gas inflow timescale. exceeding unity, favoring inflow-dominated evolution over immediate star formation. Notably, in our case, the turbulence and instability within the ISM are more likely driven by the dynamical effects of the merger itself, rather than by starburst-induced feedback, which is minimal or absent in the quiescent companion galaxy.
@@ 5.3 The Dragon Effect: AGN-Driven Negative Feedback on Companion Galaxies
While AGN and starburst activities are not detected in the Bullet, they are clearly apparent in the Host. The presence of ongoing AGN feedback in the Host galaxy is confirmed by the X-ray detection along with the presence of a broad H emission line. We thus explore the idea that the AGN in a galaxy may exert negative feedback on the companion galaxy, a possibility that has been discussed in recent studies (Davies et al., 2020; Pérez-González et al., 2024; Balashev et al., 2025; Suzuki et al., 2025). We refer to this mechanism as the Dragon Effect: a process by which star formation is effectively quenched in the companion galaxy due to AGN-driven outflow from its neighbor. This disruption can delay or suppress future star formation for extended periods, representing a novel form of external feedback in galaxy evolution and a new angle in environmental quenching at cosmic noon, where the peak of mergers, quenching, and AGN activity occurs. This scenario is particularly relevant given that the number density of AGN is known to peak in merging systems (e.g., Springel et al., 2005; Ellison et al., 2025; Euclid Collaboration et al., 2025), suggesting that such interactions—and their feedback effects—may be more common and influential than previously thought.
In a recent study, Balashev et al. (2025) report the discovery of a major merging system at redshift , where they show that radiation from a quasar hosted by one of the galaxies directly impacts the ISM of its companion. Their analysis reveals that in regions exposed to the quasar’s radiation, low-density molecular gas is significantly disrupted, leaving behind only highly compact ( number density), dense clouds that are insufficient to sustain star formation. These findings highlight a form of localized negative feedback—driven by the merger itself—that fundamentally alters the internal gas structure and likely suppresses star-forming activity. Additionally, a statistical analysis by Suzuki et al. (2025) of quasar environments at , irrespective of alignment with quasar outflow trajectories, reveals a notable deficit of Ly Emitters (LAEs) within quasar proximity regions compared to continuum-selected galaxies. The suppression is most pronounced for LAEs with high Ly equivalent widths, and both LAE and continuum galaxy number densities decline with decreasing distance to the quasar. These findings support the interpretation that quasar radiation can photoevaporate gas in low-mass halos—which typically Host LAEs—and may even impact more massive halos, thereby suppressing star formation across a range of galaxy masses.
@@ 6 Summary and Conclusion
This system thus provides the first direct observational connection between a major merger timeline and the onset of bursting and quenching phases in massive galaxies. More broadly, it offers a rare case study in how major mergers and AGN feedback can jointly regulate star formation during cosmic noon—an epoch marked by the peak of galaxy assembly, star formation, and black hole growth. Further observations of the distribution of molecular gas in such systems, along with spatially resolved investigations of their ISM properties, can shed light on the physics behind massive galaxy quenching and help determine the dominant physical mechanisms at play.
@@ Appendix A Fate of The System
where is the sum of the masses of the two galaxies (including both baryonic and dark matter components). Since the separation () is not particularly large—especially near the upper limit of —compared to the system’s size (–, the typical half-mass radius at this redshift and stellar mass; van Dokkum et al. 2010), we adopt a conservative assumption that only the stellar masses contribute to (i.e., ), neglecting the contribution of the dark matter halo. Although this approach underestimates the absolute value of the gravitational potential energy, it ensures that the point-mass approximation remains sufficiently accurate as long as the stellar mass distributions of the two galaxies have minimal overlap at their current separation.
@@
With , we find for . Given that systems with and are expected to merge rapidly (Binney & Tremaine, 1987), this implies that for the system will undergo a rapid merger. Given the current relative velocity () and the time elapsed since the impact, the relative distance and average relative velocity can be estimated for . This yields kpc and an impact velocity of km/s, values that are highly improbable even under extreme conditions, and well beyond the range of observed velocity limits (e.g., Marzke et al., 1995; Struble & Rood, 1999; Markevitch et al., 2004; Jing & Börner, 2004). Therefore, given the energy of the system, we conclude that the system is very likely to coalesce rapidly.

### v2

@@ 4.2 Stellar Kinematics
We measure the stellar kinematics in each region using pPXF (Cappellari2017) to model the absorption lines in the Y-band spectra. The stellar continuum is modeled with 32 simple stellar population templates from the EMILES library (Vazdekis2016). These templates are based on the Chabrier2003 initial mass function, BaSTI isochrones (Pietrinferni2004), and include eight stellar ages spanning from 0.15 to 14 Gyr, logarithmically spaced in steps of 0.22 dex. Four metallicity values were considered: [Z/H] = –1.5, –0.35, 0.06, and 0.4. During the fitting process, we applied multiplicative polynomials of sixth order to account for variations in the spectral shape. Note that the H emission line was fit with a velocity allowed to vary within km s-1 relative to the stellar component, to accommodate potential kinematic offsets.
@@
In conclusion, analysis of the kinematics shows that the Bullet galaxy is moving toward the observer, while the Upper Ring is gas-rich and moving away from the observer. While the Bullet and the Upper Ring appear connected in the imaging, this must be clearly a projection effect, with the Bullet likely being in the foreground of the Upper Ring.
@@ 4.3 Prospector Fit and SFHs
Similarly to what done with the photometry and explained in Sec. 3, we use prospector to model the SED of different regions, but this time including both photometric data and Y-band spectra. The model includes 29 free parameters that describe the stars, gas, and dust, as explained earlier, with additional parameters necessary to fit the spectroscopy, accounting for example for imperfect flux calibration and spectral uncertainties, as explained in detail by 2024Park. It is important to note that while dust emission is considered in our models, AGN and nebular emission are not included in our current model. For this reason, we mask the H emission lines in the spectra and do not include the H-band spectra in the fit.
The best-fit Prospector model and the observed data are presented in Fig. 4 for the Bullet and Host galaxies. We fit MOSFIRE Y-band spectra in combination with JWST and HST photometric data, extracted from the same spatial regions. Notably, the Host spectrum appears to exhibit a small excess absorption at the Ca K line (CaII Å), which is not fully accounted for by the stellar population template. This excess is also visible in Fig. 3, where the Y band spectrum is modeled using pPXF. The Ca K line has attracted growing attention in recent literature (2024Belli; 2025Liboni) because it is a resonant transition and thus can be used to detect cold gas in the galaxy reservoir and/or outflows, as an alternative diagnostic to the Na D line. The gas traced is mostly neutral, given that the ionization potential of Ca II (11.9 eV) is lower than that of hydrogen.
In line with the objectives of this study, Fig. 5 presents the SFHs of the Bullet and Host derived using two methods: (1) combined photometry and spectroscopy, and (2) photometry alone. For both methods, the photometric data were extracted from identical spatial regions, corresponding to the Keck boxes shown in Fig. 3 (distinct from those defined in Sec. 3). The results demonstrate a high degree of consistency between the two approaches, with both yielding comparable present-day SFRs (first age bin) and similar burst/quenching patterns. Noticeable discrepancy appears in the older age bins of the Host galaxy’s SFH. However, this variations do not affect the overall interpretation or conclusions of this work. Note that the associated uncertainties of the derived SFR in each age bin represent the confidence intervals around the maximum a posteriori estimates.
@@ 5.1 Quenching During the Impact Window
The kinematics of the collisional ring enables a reconstruction of the time elapsed since the impact between the two galaxies (see Sec. 4.1). We assume that the ring started expanding at the time of impact and that the expansion rate was constant (see also 2018Renaud and references therein). Accounting for observational uncertainties, the estimated collision time ranges from 47 to 96 Myr ago. This measurement is consistent with the ring detectability range ( Myr, 2018Renaud; 2023AMartinez-Delgado) and observed rings lifetime in previous studies at lower redshifts ( 50 to 300 Myr, 2011Conn; 2024Ditrani; 2012Smith). The 47 Myr lower bound corresponds to the maximum inclination and minimum rotational velocity. Recent observational (e.g., 2012Scudder; 2013Patton) and simulation studies (e.g., 2016Sparre; 2018Renaud) suggest that changes in the SFH of interacting galaxies can begin well before first contact or final coalescence (at separations kpc), sometimes as early as Myr prior to impact. Incorporating this pre-collision influence, we estimate a window of approximately 50-150 Myr over which the interaction can impact the SFH.
@@
The timeline of the impact is compared to the SFH of the Host and the Bullet in Fig. 6. As illustrated by the shaded regions in the figure, the impact window timeline (50–150 Myr) aligns closely with the onset of major changes in star formation activity—specifically, a pronounced starburst in the Host galaxy and rapid quenching in the Bullet. This temporal correspondence highlights the pivotal role of the merger in driving these evolutionary transitions. Quantitatively, the Host exhibits a star formation rate more than 25 times higher than at the onset of the impact window, while the Bullet’s star formation rate declines to nearly one-tenth of its pre-impact value, demonstrating contrasting responses by the two galaxies to the same interaction event. As discussed by (2018Renaud), variations in the SFH of collisional systems typically peak 50–150 Myr after the interaction, consistent with the rapid evolution observed at t = 0 in our system (see Fig. 5).
@@
The alignment between the timeline of quenching and that of the galaxy interaction strongly suggests the presence of a physical connection between these two processes. However, in principle it is possible that this is merely a coincidence, and the Bullet was instead quenched by an unrelated mechanism while it was also interacting with the Host. The likelihood of a massive galaxy undergoing quenching independently of an external impact can be roughly estimated. 2023Park provide number density estimates for recently quenched galaxies at redshift with stellar masses comparable to our system. They report that the number density of galaxies that quenched rapidly within the past 300 Myr at this redshift is approximately . According to 2023Weaver, based on COSMOS2020 data, the total number density of galaxies with comparable stellar mass at is about , with quiescent galaxies comprising roughly of that population. This means that if we select a massive quiescent galaxy at that is in a major merger, the probability that this system is undergoing quenching within this short timescale purely by chance—i.e., independently of the interaction with the companion—is of the order of . This low probability strengthens the case for merger-induced quenching in our system.
@@ 5.2 Alternative Process Associated with Galaxy Mergers
Quenching in merging systems is typically believed to occur due to the alteration of gas inflow, typically by rapidly consuming the gas reservoir of the galaxy during significant bursts of star formation (e.g., 2022Ellison; 2023Li). However, we rule out the possibility that quenching in the Bullet is caused by gas exhaustion based on the absence of any detectable starburst signature in the SFH (see Figs. 2 and 5). This conclusion holds even when adopting a bursty SFH model in Prospector (see Sec. 3). Another possibility is that the interaction triggers an episode of strong AGN feedback in the Bullet, which may rapidly remove the ISM gas content and shut down the star formation. In this classic scenario (e.g., 2005DiMatteo; 2008Hopkins), the black hole accretion is also typically accompanied by a substantial starburst, which we do not detect. In contrast, the Host galaxy exhibits both a significant starburst and AGN activity, indicating substantial gas inflow during the collision event. This suggests a gas-rich environment, where gas is funneled into the central regions as a result of the impact.
Several studies have proposed that galaxy interactions can induce significant turbulence within the ISM, which in turn may suppress star formation efficiency by disrupting the gravitational collapse of gas clouds (e.g., 2018Ellison; 2018vandeVoort; 2025Suess). This phenomenon is sometimes associated with so-called violent disc instability, a process that can be triggered during major mergers (e.g., 2014Dekel&Burkert; 2015Zolotov). In the context of our system, such a scenario remains plausible, particularly if the timescale for gas inflow is shorter than the timescale required for star formation. This condition corresponds to a wetness parameter333This parameter is defined as the ratio of the timescale required for star formation over gas inflow timescale. exceeding unity, favoring inflow-dominated evolution over immediate star formation. Notably, in our case, the turbulence and instability within the ISM are more likely driven by the dynamical effects of the merger itself, rather than by starburst-induced feedback, which is minimal or absent in the quiescent companion galaxy. Future investigations employing spatially resolved studies of turbulent and shocked regions will be crucial for developing a deeper understanding of this scenario.
@@ 5.3 The Dragon Effect: AGN-Driven Negative Feedback on Companion Galaxies
While AGN and starburst activities are not detected in the Bullet, they are clearly apparent in the Host. The presence of ongoing AGN feedback in the Host galaxy is confirmed by the X-ray detection along with the presence of a broad H emission line. We thus explore the idea that the AGN in a galaxy may exert negative feedback on the companion galaxy, a possibility that has been discussed in recent studies (2020Davies; 2024Perez-Gonzalez; 2025Balashev; 2025Suzuki). We refer to this mechanism as the Dragon Effect: a process by which star formation is effectively quenched in the companion galaxy due to AGN-driven outflow from its neighbor. This disruption can delay or suppress future star formation for extended periods, representing a novel form of external feedback in galaxy evolution and a new angle in environmental quenching at cosmic noon, where the peak of mergers, quenching, and AGN activity occurs. This scenario is particularly relevant given that the number density of AGN is known to peak in merging systems (e.g., 2005Springel; 2025Ellison; 2025Euclid), suggesting that such interactions—and their feedback effects—may be more common and influential than previously thought.
In a recent study, 2025Balashev report the discovery of a major merging system at redshift , where they show that radiation from a quasar hosted by one of the galaxies directly impacts the ISM of its companion. Their analysis reveals that in regions exposed to the quasar’s radiation, low-density molecular gas is significantly disrupted, leaving behind only highly compact ( number density), dense clouds that are insufficient to sustain star formation. The effect of the quasar’s intense UV radiation on molecular gas is strikingly similar to the disruptive influence of massive newborn stars on their natal clouds, fundamentally altering the internal gas structure and likely suppressing star-forming activity. Additionally, a statistical analysis by 2025Suzuki of quasar environments at , irrespective of alignment with quasar outflow trajectories, reveals a notable deficit of Ly Emitters (LAEs) within quasar proximity regions compared to continuum-selected galaxies. The suppression is most pronounced for LAEs with high Ly equivalent widths, and both LAE and continuum galaxy number densities decline with decreasing distance to the quasar. These findings support the interpretation that quasar radiation can photoevaporate gas in low-mass halos—which typically host LAEs—and may even impact more massive halos, thereby suppressing star formation across a range of galaxy masses.
@@ 6 Summary and Conclusion
This system thus provides the first direct observational connection between a major merger timeline and the onset of bursting and quenching phases in massive galaxies. More broadly, it offers a rare case study in how major mergers and AGN feedback can jointly regulate star formation during cosmic noon—an epoch marked by the peak of galaxy assembly, star formation, and black hole growth. Further observations of the distribution of molecular gas in such systems, along with spatially resolved investigations of their ISM properties, can shed light on the physics behind massive galaxy quenching and help determine the dominant physical mechanisms at play. Additionally, future multi-phase hydrodynamic simulations will be able to fully assess the Dragon Effect by quantifying its dependence on AGN power, impact parameter, and the gas structure of the companion.
7 Data Availability
Some of the data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/yxwb-6d35 (catalog DOI: 10.17909/yxwb-6d35).
@@ Appendix A Fate of The System
where is the sum of the masses of the two galaxies (including both baryonic and dark matter components). Since the separation () is not particularly large—especially near the upper limit of —compared to the system’s size (–, the typical half-mass radius at this redshift and stellar mass; 2010van_Dokkum), we adopt a conservative assumption that only the stellar masses contribute to (i.e., ), neglecting the contribution of the dark matter halo. Although this approach underestimates the absolute value of the gravitational potential energy, it ensures that the point-mass approximation remains sufficiently accurate as long as the stellar mass distributions of the two galaxies have minimal overlap at their current separation.
@@
With , we find for . Given that systems with and are expected to merge rapidly (1987Binney), this implies that for the system will undergo a rapid merger. For the current relative velocity is , the current separation is kpc and the mean relative speed (computed as current separation divided by time elapsed since the collision) is . A smaller value of and larger values of and would be found for . The inferred implies a relative speed at the time of collision far exceeding —a value highly implausible even under extreme conditions and well beyond observed limits (e.g., 1995Marzke; 1999Struble; 2004Markevitch; 2004Jing). Therefore, given the energy of the system, we conclude that the system is very likely to coalesce rapidly.